Show that this rearranges to q/m=v/BR


1
Expert's answer
2022-06-07T08:11:51-0400

"F_1=qvB,"

"F_2=\frac{mv^2}r,"

"F_1=F_2,"

"qB=\frac{mv}r,\implies"

"\frac qm=\frac v{Br}."